# Improving Graph Neural Networks by Learning Continuous Edge Directions

### v1

@@ Abstract
Graph Neural Networks (GNNs) traditionally employ a message-passing mechanism that resembles diffusion over undirected graphs, which often leads to homogenization of node features and reduced discriminative power in tasks such as node classification. Our key insight for addressing this limitation is to assign fuzzy edge directions—that can vary continuously from node $i$ pointing to node $j$ to vice versa—to the edges of a graph so that features can preferentially flow in one direction between nodes to enable long-range information transmission across the graph. We also introduce a novel complex-valued Laplacian for directed graphs with fuzzy edges where the real and imaginary parts represent information flow in opposite directions. Using this Laplacian, we propose a general framework, called Continuous Edge Direction (CoED) GNN, for learning on graphs with fuzzy edges and prove its expressivity limits using a generalization of the Weisfeiler-Leman (WL) graph isomorphism test for directed graphs with fuzzy edges. Our architecture aggregates neighbor features scaled by the learned edge directions and processes the aggregated messages from in-neighbors and out-neighbors separately alongside the self-features of the nodes. Since continuous edge directions are differentiable, they can be learned jointly with the GNN weights via gradient-based optimization. CoED GNN is particularly well-suited for graph ensemble data where the graph structure remains fixed but multiple realizations of node features are available, such as in gene regulatory networks, web connectivity graphs, and power grids. We demonstrate through extensive experiments on both synthetic and real datasets that learning continuous edge directions significantly improves performance both for undirected and directed graphs compared with existing methods.
00footnotetext: Our code is available at https://github.com/hormoz-lab/coed-gnn.
@@ 1 Introduction
Graph Neural Networks (GNNs) have emerged as a powerful tool for learning from data that is structured as graphs, with applications ranging from social network analysis to molecular chemistry (Kipf & Welling, 2017; Zhou et al., 2020; Gilmer et al., 2017). GNNs typically employ a message passing mechanism where nodes aggregate and then transform feature information from their neighbors at each layer, enabling them to learn node representations that capture both local and global graph structures. When the graph is undirected, the aggregation of node features mimics a diffusion process. Each node’s representation becomes the averaged features of its immediate neighbors, leading to a homogenization of information across the graph. As depth increases, this diffusion of information culminates in a uniform state where node representations converge towards a constant value across all the nodes, which severely limits the discriminative power of GNNs, especially in tasks such as node classification (Rusch et al., 2023a; Oono & Suzuki, 2020; Cai & Wang, 2020; Li et al., 2018a; Keriven, 2022; Chen et al., 2020a; Wu et al., 2022; 2024).
Figure 1: (a) When edges are undirected information diffuses across the graph and long range transmission of information between nodes 1 and 2 is not possible. (b) Once the optimal edge directions are learned, information can flow directly from node 1 to node 2.
Our key insight for improving the performance of GNNs is to alter the nature of information transmission between nodes from diffusion to flow. To do so, we add directions to the edges of a graph so that features can be propagated from node $v_{i}$ to its neighbor node $v_{j}$ without a reciprocal propagation of information from node $v_{j}$ to node $v_{i}$. Unlike diffusion, where information uniformly spreads across available paths, flow is directional and preserves the propagation of information across longer distances within a graph, as illustrated in Figure 1. In general, the optimal information propagation could require edges whose directions fall anywhere in the continuum of node $v_{i}$ pointing to node $v_{j}$ to vice versa.
@@
Next, we introduce the Continuous Edge Direction (CoED) GNN architecture. At each layer, a node’s neighbors’ features are scaled by the directions of their connecting edges and aggregated. This aggregation is performed separately for incoming and outgoing edges, following Rossi et al. (2024), resulting in distinct features for incoming and outgoing messages. Practically, this is implemented by applying the fuzzy Laplacian to the node features, where the real and imaginary parts correspond to the features aggregated from incoming neighbors and outgoing neighbors, respectively. These aggregated features are then affine transformed using learnable weights and combined with the node’s own transformed features. A nonlinear activation function is applied to obtain the updated node features. This process is repeated for each layer. The continuous edge directions have the added benefit that they are differentiable. During training, both the edge directions and weight matrices are learned simultaneously using gradient-based optimization to improve the learning objective.
@@
Importantly, our approach is fundamentally different from methods such as Graph Attentions Network (GAT) (Veličković et al., 2018) or graph transformers (Dwivedi & Bresson, 2021; Rampášek et al., 2022) that learn attention coefficients to assign weights to each edge of the graph based on the features (and potentially the positional encoding) of the nodes connected by that edge. While such an attention mechanism can capture asymmetric relationships by computing direction-specific attention weights based on node features, they do not learn edge directions as independent parameters. In these models, the attention coefficient from node $v_{i}$ to node $v_{j}$ are functions of the features of $v_{i}$ and $v_{j}$, and will change if node features change. In contrast, our approach introduces continuous edge directions as learnable parameters that are optimized end-to-end, independent of the node features. As we demonstrate empirically below, by directly learning edge directions, our method goes beyond what can be achieved through an attention mechanism, enabling long-range information flow on graphs and improved performance.
@@
End-to-end learning of edge directions is most effective on graph ensemble data, where the graph structure remains fixed but multiple realizations of node features and targets (such as node labels) are available. This effectiveness arises from the ability to optimize information flow across all edges simultaneously without a need to mask parts of the graph for training and testing. Instead, training and testing splits are based on different feature realizations rather than on subsets of the graph. Graph ensemble data are increasingly common across various domains. For example, in biology, gene-regulatory networks are constant directed graphs where nodes represent genes and edges represent gene-gene interactions, while node features like gene expression levels vary across different cells. Similarly, in web connectivity, the network of websites remains relatively static, but traffic patterns change over time, providing different node feature sets on the same underlying graph. In power grids, the network of electrical components is fixed, while the steady-state operating points of these components vary under different conditions, yielding multiple observations on the same graph. In all these cases, a fixed graph is paired with numerous feature variations. By applying CoED GNN to these scenarios, we demonstrate that learning edge directions significantly improves performance for both directed and undirected graphs. The main contributions of this paper are the following:
@@
  • Using extensive experiments, we empirically show that learning edge directions significantly improves performance by applying CoED GNN to both synthetic and real graph ensemble data.
@@ 2 Preliminaries
In GNNs, node features $\mathbf{F}$ are processed iteratively through a message-passing mechanism that leverages the structural information of the graph $\mathcal{G}$. This process involves two main steps at each layer $l$:
@@ 3.2 Fuzzy Graph Laplacian
$(\mathbf{L}_{F})_{ij}=\begin{cases}0&\text{if }A_{ij}=A_{ji}=0\\ \exp(i\theta_{ij})&\text{otherwise}\end{cases}$ (1)
Since $\theta_{ij}$ and $\theta_{ji}$ are related by $\theta_{ji}=\pi/2-\theta_{ij}$, it follows that $\mathbf{L}_{F}=i\mathbf{L}_{F}^{\dagger}$, where $\dagger$ is the conjugate transpose. $\text{Re}[{\mathbf{L}_{F}}]$ thus encodes all $i\leftarrow j$ edges scaled by $\cos{\theta_{ij}}$ and $\text{Im}[{\mathbf{L}_{F}}]$ all $i\rightarrow j$ edges scaled by $\sin{\theta_{ij}}$. In Appendix D, we show the fuzzy graph Laplacian admits orthogonal eigenvectors with eigenvalues of the form $a+ia$ with $a\in\mathbb{R}$. Therefore, the eigenvectors of our Laplacian provide positional encodings that are informed by the directions of the edges in addition to their connectivities, as shown in Appendix C. In addition, we prove in Appendix E that a message-passing GNN whose aggregation step is done using the fuzzy Laplacian is as expressive as an extension of the Weisfeiler-Leman (WL) graph isomorphism test to directed graphs with fuzzy edges.
An alternative form of Laplacian for directed graphs is the magnetic Laplacian (Zhang et al., 2021; He et al., 2022), which is also complex-valued. For directed graphs with fuzzy edges, the magnetic Laplacian is not as expressive as the Laplacian proposed above. We provide a proof in Appendix F. Briefly, the magnetic Laplacian produces linear combinations of in- and out- neighbor messages at each node as both the real and imaginary parts of the aggregated features. In principle, GNNs should be able to disentangle these linear combinations to recover the in- and out- neighbor messages. However, the linear combinations depend on the local neighborhood of each node which is distinct from one node to another, whereas GNN parameters are shared across all nodes. Therefore, in general, a GNN using the magnetic Laplacian loses the ability to disentangle the in- and out- neighbor messages at each node and thus has lower expressivity. The Laplacian proposed above does not suffer from this limitation since by construction the real and imaginary values directly correspond to the in- and out- neighbor aggreagated messages, respectively.
@@ 3.3 Model Architecture: Continuous Edge Direction (CoED) GNN
To ensure maximum expressivity, a message-passing mechanism on a directed graph should, for each node, separately aggregate the features of the in-neighbors and the out-neighbors, and independently process each of the aggregated features and the self-features to obtain an updated feature for each node. To this end, we define in- and out- edge weight matrices as $\mathbf{A_{\leftarrow}}=\text{Re}[{\mathbf{L}_{F}}]$ and $\mathbf{A_{\rightarrow}}=\text{Im}[{\mathbf{L}_{F}}]$, respectively. We compute the in- and out- degree matrices as $\mathbf{D_{\leftarrow}}=\mathrm{diag}\left(\mathbf{A_{\leftarrow}}\mathbf{1}\right)$ and $\mathbf{D_{\rightarrow}}=\mathrm{diag}\left(\mathbf{A_{\rightarrow}}\mathbf{1}\right)$. Following Rossi et al. (2024), but extending it to graphs with continuous edge directions, we define in- and out- fuzzy propagation matrices as
$\begin{split}\mathbf{P}_{\leftarrow}=\mathbf{D}^{-1/2}_{\leftarrow}\mathbf{A}_{\leftarrow}\mathbf{D}^{-1/2}_{\rightarrow}\\ \mathbf{P}_{\rightarrow}=\mathbf{D}^{-1/2}_{\rightarrow}\mathbf{A}_{\rightarrow}\mathbf{D}^{-1/2}_{\leftarrow}\end{split}$ (2)
@@
$\begin{split}\mathbf{m}^{(l)}_{\leftarrow}=\mathbf{P}_{\leftarrow}\mathbf{F}^{(l-1)}\\ \mathbf{m}^{(l)}_{\rightarrow}=\mathbf{P}_{\rightarrow}\mathbf{F}^{(l-1)}\end{split}$ (3)
@@
The features at the final layer are then transformed using a linear layer to obtain the output for a specific learning task, e.g. node classification or node regression. We use end-to-end gradient-based optimization to iteratively update both the phase matrix $\Theta$ and the GNN parameters $\mathbf{W}_{\text{self}/\leftarrow/\rightarrow}^{(l)}$ and $\mathbf{B}^{(l)}$ at each layer, as illustrated in Figure 2. We allow for the option to learn a different set of edge directions at each layer, $\Theta^{(l)}$, just as we have distinct GNN parameters at each layer.
@@ 4 Related Work
Extending GNNs to directed graphs has been addressed through various methods. GatedGNN (Li et al., 2016) processed messages from out-neighbors in directed graphs. Some works constructed symmetric matrices from directed adjacency matrices and their transposes to build standard Laplacians (Tong et al., 2020b; Kipf & Welling, 2017), while others (Ma et al., 2019; Tong et al., 2020a) developed Laplacians based on random walks and PageRank (Duhan et al., 2009). MagNet (Zhang et al., 2021) utilized the magnetic Laplacian to represent directed messages, a technique also applied in adapting transformers to directed graphs (Geisler et al., 2023). FLODE (Maskey et al., 2023) employed asymmetrically normalized adjacency matrices within a neural ODE framework. DirGNN (Rossi et al., 2024) separately processed the messages from in-neighbors and out-neighbors using asymmetrically normalized adjacency matrices, improving node classification on heterophilic graphs. A similar strategy was used in Koke & Cremers (2024), replacing the adjacency matrices with filters representing Faber polynomials. Recent graph PDE-based models (Eliasof et al., 2024; Zhao et al., 2023) introduced an advection term to model directional feature propagation alongside diffusion, assigning edge weights based on computed velocities between nodes, akin to attention coefficients in GAT.
@@ 5 Experiments
5.1 Node classification
@@
We first benchmarked our method for node classification task on five graphs with varying properties without simultaneously learning the edge directions. While this is not the main goal of our paper, we include this analysis nonetheless to highlight the advantage of our Laplacian over the magnetic Laplacian as well as the benefit of processing self, in-neighbor aggregated, and out-neighbor aggregated features separately. Importantly, we do not learn edge directions in this case, and hence the phase value is either 00 or $\pi/2$ for directed graphs and $\pi/4$ for undirected graphs. For comparison, we include the classical models: GCN (Kipf & Welling, 2017), SAGE (Hamilton et al., 2017), GAT (Veličković et al., 2018); heterophily-specific model, GGCN (Yan et al., 2021); directionality-aware models: MagNet (Zhang et al., 2021), FLODE (Maskey et al., 2023), DirGNN (Rossi et al., 2024); and a model that learns geometric structure of graph, Sheaf (Bodnar et al., 2022). We also include MLP to highlight the effect of solely processing the nodes’ self-features without aggregating features across the graph.
@@
As shown in Table 1, CoED demonstrates competitive performance across all five datasets, ranking within the top three in terms of test accuracy. While all models exhibit comparable results on the Cora graph—which is undirected and homophilic—their performances differ significantly on the directed, heterophilic graphs. The classical models developed for undirected graphs particularly struggle on these datasets, with the exception of SAGE, which achieves over 80% accuracy on Texas and Wisconsin. This is because processing only the node’s own features yields good performance, as evidenced by the MLP’s results. In contrast, for the Squirrel and Chameleon datasets, processing directed messages along only one direction is crucial for good performance. Only FLODE, DirGNN, and CoED exhibit strong results on these datasets when configured accordingly. Specifically, for CoED, we introduce the $\alpha$ hyperparameter as in Rossi et al. (2024) to weigh the directional messages post aggregation and make the transformation of self-features optional. The selected hyperparameters modify Equation 4 to $\sigma(\mathbf{m}^{(l)}_{\leftarrow}\mathbf{W}_{\leftarrow}^{(l)}+\mathbf{B}^{(l)})$ for these datasets.
@@
Our results highlight the advantage of the fuzzy Laplacian over the magnetic Laplacian. The magnetic Laplacian does not process the aggregated messages from out-neighbors and in-neighbors separately. Instead, it combines them into both the real and imaginary components of the aggregated feature vector, thus losing the opportunity to process the two separately. Moreover, during the Laplacian convolution, the directed messages further mix with self-features encoded in the real component. This results in poor performance by MagNet on Squirrel and Chameleon. Sheaf also suffers on these datasets despite expanding the feature dimensions via an object called a stalk, because the sheaf Laplacian is constrained to be symmetric, thereby losing the ability to process directed messages. FLODE and DirGNN, however, perform poorly on Texas and Wisconsin because they focus exclusively on directed messages and do not process the node’s own features separately. Taken together, our benchmarking demonstrates that CoED’s ability to effectively process self-features and separately aggregate in-neighbor and out-neighbor messages using our fuzzy Laplacian enables it to achieve competitive performance across diverse datasets.
Texas Wisconsin Squirrel Chameleon Cora
Hom. level 0.11 0.21 0.22 0.23 0.81

### v2

@@ Abstract
Graph Neural Networks (GNNs) traditionally employ a message-passing mechanism that resembles diffusion over undirected graphs, which often leads to homogenization of node features and reduced discriminative power in tasks such as node classification. Our key insight for addressing this limitation is to assign fuzzy edge directions—that can vary continuously from node $i$ pointing to node $j$ to vice versa—to the edges of a graph so that features can preferentially flow in one direction between nodes to enable long-range information transmission across the graph. We also introduce a novel complex-valued Laplacian for directed graphs with fuzzy edges where the real and imaginary parts represent information flow in opposite directions. Using this Laplacian, we propose a general framework, called Continuous Edge Direction (CoED) GNN, for learning on graphs with fuzzy edges and prove its expressivity limits using a generalization of the Weisfeiler-Leman (WL) graph isomorphism test for directed graphs with fuzzy edges. Our architecture aggregates neighbor features scaled by the learned edge directions and processes the aggregated messages from in-neighbors and out-neighbors separately alongside the self-features of the nodes. Since continuous edge directions are differentiable, they can be learned jointly with the GNN weights via gradient-based optimization. CoED GNN is particularly well-suited for graph ensemble data where the graph structure remains fixed but multiple realizations of node features are available, such as in gene regulatory networks, web connectivity graphs, and power grids. We demonstrate through extensive experiments on both synthetic and real graph ensemble datasets that learning continuous edge directions significantly improves performance both for undirected and directed graphs compared with existing methods. Our code is available on GitHub.
[…]
@@ 1 Introduction
Graph Neural Networks (GNNs) have emerged as a powerful tool for learning from data that is structured as graphs, with applications ranging from social network analysis to molecular chemistry (Kipf & Welling, 2017; Zhou et al., 2020; Gilmer et al., 2017). GNNs typically employ a message passing mechanism where nodes aggregate and then transform feature information from their neighbors at each layer, enabling them to learn node representations that capture both local and global graph structures. When the graph is undirected, the aggregation of node features mimics a diffusion process. Each node’s representation becomes the averaged features of its immediate neighbors, leading to a homogenization of information across the graph. As depth increases, this diffusion of information culminates in a uniform state where node representations converge towards a constant value across all nodes, which severely limits the discriminative power of GNNs, especially in tasks such as node classification (Rusch et al., 2023a; Oono & Suzuki, 2020; Cai & Wang, 2020; Li et al., 2018a; Keriven, 2022; Chen et al., 2020a; Wu et al., 2023; 2024).
Figure 1: (a) When edges are undirected, information diffuses across the graph and long-range transmission of information between nodes 1 and 2 is not possible. (b) Once the optimal edge directions are learned, information can flow directly from node 1 to node 2.
Our key insight for improving the performance of GNNs is to alter the nature of information transmission between nodes from diffusion to flow. To do so, we add directions to the edges of a graph so that features can be propagated from node $v_{i}$ to its neighbor node $v_{j}$ without reciprocal propagation of information from node $v_{j}$ to node $v_{i}$. Unlike diffusion, where information uniformly spreads across available paths, flow is directional and preserves the propagation of information across longer distances within a graph, as illustrated in Figure 1. In general, the optimal information propagation could require edges whose directions fall anywhere in the continuum of node $v_{i}$ pointing to node $v_{j}$ to vice versa.
@@
Next, we introduce the Continuous Edge Direction (CoED) GNN architecture. At each layer, a node’s neighbors’ features are scaled by the directions of their connecting edges and aggregated. This aggregation is performed separately for incoming and outgoing edges, following Rossi et al. (2024), resulting in distinct features for incoming and outgoing messages. This is implemented by applying the fuzzy Laplacian to the node features, where the real and imaginary parts correspond to the features aggregated from incoming neighbors and outgoing neighbors, respectively. These aggregated features are then affine transformed using learnable weights and combined with the node’s own transformed features. A nonlinear activation function is applied to obtain the updated node features. This process is repeated for each layer. The continuous edge directions have the added benefit that they are differentiable. During training, both the edge directions and weight matrices are learned simultaneously using gradient-based optimization to improve the learning objective.
@@
Importantly, our approach is fundamentally different from methods such as Graph Attentions Network (GAT) (Veličković et al., 2018) or graph transformers (Dwivedi & Bresson, 2021; Rampášek et al., 2022) that learn attention coefficients to assign weights to each edge of the graph based on the features (and potentially the positional encoding) of the nodes connected by that edge. While the attention mechanism can capture asymmetric relationships by computing direction-specific attention weights based on node features, they do not learn edge directions as independent parameters. Instead, the attention coefficients from node $v_{i}$ to node $v_{j}$ are functions of the features of $v_{i}$ and $v_{j}$, and will change if node features change. In contrast, our approach introduces continuous edge directions as learnable parameters that are optimized end-to-end, independent of the node features.
@@
Learning edge directions can make CoED susceptible to overfitting, especially when the same graph is used for both training and testing, as in standard node classification tasks. In such cases, edge directions may be optimized for the training nodes at the expense of effective information flow to the test nodes that are withheld during training. In contrast, CoED is very effective on graph ensemble data, where the graph structure remains fixed but multiple realizations of node features and targets (such as node labels) are available. This allows optimization of information flow across all edges simultaneously without masking parts of the graph for training and testing. Instead, training and testing splits are based on different feature realizations rather than on subsets of the graph. Graph ensemble data are increasingly common across various domains. For example, in biology, gene-regulatory networks are constant directed graphs where nodes represent genes and edges represent gene-gene interactions, while node features like gene expression levels vary across different cells. Similarly, in web connectivity, the network of websites remains relatively static, but traffic patterns change over time, providing different node feature sets on the same underlying graph. In power grids, the network of electrical components is fixed, while the steady-state operating points of these components vary under different conditions, yielding multiple observations on the same graph. In all these cases, a fixed graph is paired with numerous feature variations. By applying CoED GNN to these scenarios, we demonstrate that learning edge directions significantly improves performance for both directed and undirected graphs. The main contributions of this paper are the following:
@@
  • Using extensive experiments, we show that learning edge directions significantly improves performance by applying CoED GNN to both synthetic and real graph ensemble data.
@@ 2 Preliminaries
GNNs iteratively processes node features $\mathbf{F}$ via message-passing mechanism that leverages the structural information of the graph $\mathcal{G}$. This process involves two main steps at each layer $l$:
@@ 3.2 Fuzzy Graph Laplacian
$(\mathbf{L}_{F})_{ij}=\begin{cases}0&\text{if }A_{ij}=A_{ji}=0\\ \exp(\mathrm{i}\theta_{ij})&\text{otherwise}\end{cases}$ (1)
Since $\theta_{ij}$ and $\theta_{ji}$ are related by $\theta_{ji}=\pi/2-\theta_{ij}$, it follows that $\mathbf{L}_{F}=\mathrm{i}\mathbf{L}_{F}^{\dagger}$, where $\dagger$ is the conjugate transpose. $\text{Re}[{\mathbf{L}_{F}}]$ thus encodes all $i\leftarrow j$ edges scaled by $\cos{\theta_{ij}}$ and $\text{Im}[{\mathbf{L}_{F}}]$ all $i\rightarrow j$ edges scaled by $\sin{\theta_{ij}}$. In Appendix D, we show the fuzzy graph Laplacian admits orthogonal eigenvectors with eigenvalues of the form $a+\mathrm{i}a$ with $a\in\mathbb{R}$. Therefore, the eigenvectors of our Laplacian provide positional encodings that are informed by the directions of the edges in addition to their connectivities. We provide the visualizations of the eigenvectors in Appendix C. A key implication of our Laplacian for GNN architectures is stated in the following theorem.
Theorem 1.
A message-passing GNN whose aggregation step is performed using the fuzzy graph Laplacian is as expressive as the weak form of the Weisfeiler-Leman (WL) graph isomorphism test for directed graphs with fuzzy edges.
We prove this theorem in Appendix E. The most commonly used Laplacian for directed graphs is the magnetic Laplacian (Shubin, 1994; Furutani et al., 2020; Zhang et al., 2021; He et al., 2022a), which is also complex-valued. For directed graphs with fuzzy edges, the magnetic Laplacian is not as expressive as the Laplacian proposed above. We provide a proof in Appendix F. Briefly, the real and imaginary parts of the aggregated features produced by the magnetic Laplacian are both linear combinations of in- and out- neighbor messages. In principle, GNNs should be able to disentangle these linear combinations to recover the in- and out- neighbor messages. However, the linear combinations depend on the local neighborhood of each node which is distinct from one node to another, whereas GNN parameters are shared across all nodes. Therefore, in general, a GNN using the magnetic Laplacian loses the ability to disentangle the in- and out- neighbor messages at each node and thus has lower expressivity. Our Laplacian does not suffer from this limitation since by construction the real and imaginary values uniquely correspond to the in- and out- neighbor aggregated messages, respectively.
@@ 3.3 Model Architecture: Continuous Edge Direction (CoED) GNN
To ensure maximum expressivity, a message-passing mechanism on a directed graph should, for each node, separately aggregate the features of the in-neighbors and the out-neighbors, and independently process the two types of aggregated features and the self-features to obtain an updated feature for each node. To this end, we define in- and out- edge weight matrices as $\mathbf{A_{\leftarrow}}=\text{Re}[{\mathbf{L}_{F}}]$ and $\mathbf{A_{\rightarrow}}=\text{Im}[{\mathbf{L}_{F}}]$, respectively. We compute the in- and out- degree matrices as $\mathbf{D_{\leftarrow}}=\mathrm{diag}\left(\mathbf{A_{\leftarrow}}\mathbf{1}\right)$ and $\mathbf{D_{\rightarrow}}=\mathrm{diag}\left(\mathbf{A_{\rightarrow}}\mathbf{1}\right)$, respectively. Following Rossi et al. (2024), but extending it to graphs with continuous edge directions, we define in- and out- fuzzy propagation matrices as
$\mathbf{P}_{\leftarrow}=\mathbf{D}^{-1/2}_{\leftarrow}\mathbf{A}_{\leftarrow}\mathbf{D}^{-1/2}_{\rightarrow}\ \ \ \text{and}\ \ \ \mathbf{P}_{\rightarrow}=\mathbf{D}^{-1/2}_{\rightarrow}\mathbf{A}_{\rightarrow}\mathbf{D}^{-1/2}_{\leftarrow}$ (2)
@@
$\mathbf{m}^{(l)}_{\leftarrow}=\mathbf{P}_{\leftarrow}\mathbf{F}^{(l-1)}\ \ \ \text{and}\ \ \ \mathbf{m}^{(l)}_{\rightarrow}=\mathbf{P}_{\rightarrow}\mathbf{F}^{(l-1)}$ (3)
@@
The features at the final layer are then transformed using a linear layer to obtain the output for a specific learning task. We use end-to-end gradient-based optimization to iteratively update both the phase matrix $\Theta$ and the GNN parameters $\mathbf{W}_{\text{self}/\leftarrow/\rightarrow}^{(l)}$ and $\mathbf{B}^{(l)}$ at each layer, as illustrated in Figure 2. We allow for the option to learn a different set of edge directions at each layer, $\Theta^{(l)}$, just as we have distinct GNN parameters at each layer. We provide a runtime analysis of CoED against other GNNs in Appendix G.
@@ 4 Related Work
Extending GNNs to directed graphs has been addressed through various methods. GatedGNN (Li et al., 2016) processed messages from out-neighbors in directed graphs. Some works constructed symmetric matrices from directed adjacency matrices and their transposes to build standard Laplacians (Tong et al., 2020b; Kipf & Welling, 2017), while others (Ma et al., 2019; Tong et al., 2020a) developed Laplacians based on random walks and PageRank (Duhan et al., 2009). MagNet (Zhang et al., 2021) utilized the magnetic Laplacian to represent directed messages—a technique that has also been used to adapt transformers to directed graphs (Geisler et al., 2023) and to visualize them (Fanuel et al., 2018). FLODE (Maskey et al., 2023) employed asymmetrically normalized adjacency matrices within a neural ODE framework. DirGNN (Rossi et al., 2024) separately processed the messages from in-neighbors and out-neighbors using asymmetrically normalized adjacency matrices, improving node classification on heterophilic graphs. A similar strategy was used in Koke & Cremers (2024), replacing the adjacency matrices with filters representing Faber polynomials. Recent graph PDE-based models (Eliasof et al., 2024; Zhao et al., 2023) introduced an advection term to model directional feature propagation alongside diffusion, assigning edge weights based on computed velocities between nodes, akin to attention coefficients in GAT. Finally, our approach is conceptually related to He et al. (2022b) where imbalance of incoming and outgoing messages across subsets of nodes is used to learn node embeddings for clustering.
@@ 5 Experiments
5.1 Node classification without edge direction learning
@@
While node classification is not the primary focus of our paper, we benchmarked our method on eleven standard datasets—including both undirected and directed graphs covering a wide range of sizes and homophily levels—to highlight the advantage of our Laplacian over alternative forms of Laplacians for directed graphs as well as the benefit of processing self, in-neighbor, and out-neighbor features separately. Importantly, we do not learn edge directions in this case, and hence the phase value is either 00 or $\pi/2$ for directed graphs and $\pi/4$ for undirected graphs. For comparison, we include the classical models: GCN (Kipf & Welling, 2017), SAGE (Hamilton et al., 2017), GAT (Veličković et al., 2018); heterophily-specific model, GGCN (Yan et al., 2021); directionality-aware models: MagNet (Zhang et al., 2021), FLODE (Maskey et al., 2023), DirGNN (Rossi et al., 2024); and a model that learns geometric structure of graph, Sheaf (Bodnar et al., 2022). We also include a model based on a Laplacian for directed graphs constructed from the transition matrix of the graph by Chung (2005), and Cooperative GNNs (Finkelshtein et al., 2023), which classify a node as broadcasting, listening, both, or neither based on its own and its neighbors’ features. Finally, we also include MLP to highlight the effect of solely processing the nodes’ self-features without aggregating features across the graph. Further details of the datasets and hyperparameters are provided in Appendix A.1
@@
As shown in Table 1, CoED demonstrates competitive performance across all eleven datasets, ranking within the top three in terms of test accuracy for most. While all models exhibit comparable results on Cora and Citeseer—which are undirected and homophilic—their performances differ significantly on the directed, heterophilic graphs. The classical models developed for undirected graphs particularly struggle on these datasets, with the exception of SAGE. This is because processing only the node’s own features yields good performance, as evidenced by the MLP’s results. In contrast, for the Squirrel and Chameleon datasets, processing directed messages along only one direction is crucial for good performance. Only FLODE, DirGNN, and CoED exhibit strong results on these datasets when configured accordingly. Specifically, for CoED, we introduce the $\alpha$ hyperparameter as in Rossi et al. (2024) to weigh the directional messages post aggregation, replacing $\mathbf{m}^{(l)}_{\leftarrow}\mathbf{W}_{\leftarrow}^{(l)}+\mathbf{m}^{(l)}_{\rightarrow}\mathbf{W}_{\rightarrow}^{(l)}$ in Equation 4 with $\alpha\mathbf{m}^{(l)}_{\leftarrow}\mathbf{W}_{\leftarrow}^{(l)}+(1-\alpha)\mathbf{m}^{(l)}_{\rightarrow}\mathbf{W}_{\rightarrow}^{(l)}$. In addition, we make the transformation of self-features optional.
@@
Importantly, our results highlight the advantage of the fuzzy Laplacian over the magnetic Laplacian and the Chung Laplacian. In particular, the magnetic Laplacian does not process the aggregated messages from out-neighbors and in-neighbors separately. Instead, it combines them into both the real and imaginary components of the aggregated feature vector, thus losing the opportunity to process the two separately. Moreover, during the Laplacian convolution, the directed messages further mix with self-features encoded in the real component. This results in poor performance by MagNet on directed graphs. The real-world directed graphs are often not strongly connected and thus the corresponding transition matrices’ top left singular vectors are not as informative as in the undirected cases. Since the Chung Laplacian is constructed from this singular vector, it shows relatively poor performance on such datasets as Roman-Empire or Wisconsin compared to the other directed models and instead delivers better results on datasets with undirected graphs. Sheaf also suffers on the datasets with directed graphs despite expanding the feature dimensions via an object called a stalk, because the sheaf Laplacian is constrained to be symmetric, thereby losing the ability to process directed messages. Taken together, our benchmarking demonstrates that CoED’s ability to effectively process self-features and separately aggregate in-neighbor and out-neighbor messages using our fuzzy Laplacian enables it to achieve competitive performance across diverse datasets.
Roman-Empire SNAP-Patents Texas Wisconsin Arxiv-Year Squirrel Chameleon Citeseer Computers Photo Cora
Hom. level 0.05 0.07 0.11 0.21 0.22 0.22 0.23 0.74 0.78 0.81 0.81